OpenStudy (anonymous):

given that 1-i(square root(3)) is one of the cube roots of x, find x and then find its other cube roots

OpenStudy (e.mccormick):

So, \(1-i\sqrt{3}= \sqrt[3]{ x}\) is given. Where would you go from there?

OpenStudy (anonymous):

would I cube both sides?

OpenStudy (e.mccormick):

I think that is a really good apprach.

OpenStudy (anonymous):

so x=1-i3sqrt{3}?

OpenStudy (e.mccormick):

hmm.... I got something I did not expect. Going to need to check my math.

OpenStudy (e.mccormick):

Ah, canceled something wrong... I got \(1-3i\sqrt{3}+9i-i(3)^{\frac{3}{2}}\)

OpenStudy (e.mccormick):

first time I got 8 and I was like... ummm.... something went wrong. LOL.

OpenStudy (anonymous):

Ah, I know what I did wrong to get what I got previously, I forgot to foil there

OpenStudy (anonymous):

haha

OpenStudy (e.mccormick):

OK... now... here is the funny thing about this. If that is pone of the cube roots.... what does that mean the other cube roots are?

OpenStudy (anonymous):

Wait doesnt that mean all the cube roots are the same...? or am I just looking at this wrong? O.o

OpenStudy (e.mccormick):

Yep. That was what I instantly thought too. So I would go with that.

OpenStudy (anonymous):

Alright! Thanks for all your help! :D

OpenStudy (e.mccormick):

Also, beacause cube roots do retain signs, I am not worried about some possiblity of a \(\pm\sqrt{x}\) creeping in there. have fun!